# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Registration

### v1

@@ Abstract
Diffeomorphic Image Registration is a critical part of the analysis in various imaging modalities and downstream tasks like image translation, segmentation, and atlas building. Registration algorithms based on optimization have stood the test of time in terms of accuracy, reliability, and robustness across a wide spectrum of modalities and acquisition settings. However, these algorithms converge slowly, are prohibitively expensive to run, and their usage requires a steep learning curve, limiting their scalability to larger clinical and scientific studies. In this paper, we develop multi-scale Adaptive Riemannian Optimization algorithms for diffeomorphic image registration. We demonstrate compelling improvements on image registration across a spectrum of modalities and anatomies by measuring structural and landmark overlap of the registered image volumes. Our proposed framework leads to a consistent improvement in performance, and from 300$\times$ up to 2000$\times$ speedup over existing algorithms. Our modular library design makes it easy to use and allows customization via user-defined cost functions.
Keywords: image registration, image matching, image alignment, diffeomorphisms, multi-scale optimization, scalability, MRI, computed tomography, microscopy
1 Main
Deformable Image Registration is one of the most ubiquitous tasks in image analysis. It refers to the non-linear and local (hence deformable) alignment of two or more images into a common coordinate system. Depending on the problem and modality, the images can be sourced from different subjects or events, modalities, and timepoints. Image registration is routinely used in neuroimaging  1, 2, 3, 4, 5, cardiac imaging  6, 7, 8, lung imaging  9, 10, 11, microscopy and histology  12, 13, 14, 15, 16 to name a few biomedical applications. In neuroimaging, inter-subject registration is used to align structural regions for automatic segmentation, or to construct an anatomical template (atlas) for anomaly detection or deviations from a healthy population. In lung imaging, registration is used to understand the dynamics of the lung deformation during inspiration-expiration, or to track lesions over temporally spaced breathhold scans. Image registration is used in microscopy 12, 13, 16 to compensate for the large deformations that occur between staining rounds, and stitching misaligned 2D histology slides to generate a 3D volume. We note that registration is often used beyond biomedical applications; in planetary image alignment  17, satellite imagery and remote sensing  18, 19, 20, robotics  21, 22, 23, and astronomy  24, 25, 26 and traditional computer vision applications like optical flow 27, 28, 29, 30. In this paper, we focus on image registration for biomedical applications, including microscopy, Magnetic Resonance Imaging (MRI), and Computed Tomography (CT) imaging, but our method is more generally applicable to other imaging as well.
Image registration methods are typically divided into two categories – optimization-based and learning-based. Optimization-based methods focus on mathematically formulating registration as a variational optimization problem. This involves selecting a dissimilarity function between the reference and warped image, the family of deformation fields over which to optimize, and the optimization algorithm to use. In the literature, the reference and warped images are typically called fixed and moving images respectively. Diffeomorphisms are of special interest as a family of deformations, which are invertible transformations such that both the transform and its inverse are differentiable. Some of the earliest approaches considered models for small deformations  31, 32, 33, 34, 35. Other approaches perform gradient based optimization on the variational objective function  36, 37, 38, 39, modelling diffeomorphisms as solutions of a differential equation with a time-dependent velocity field 40, followed up with geodesic formulations 41, 42, 43, and direct integration of the velocity fields using gradient descent 44, 45. These methods focus on the representation choice and optimization technique. An orthogonal problem in image registration is the lack of discriminative features in medical images which are noisy and contain artifacts, making intensity based registration susceptible to local minima and slow convergence. To overcome these problems, learning based methods train a deep neural network that inputs the intensity images and predicts the deformation directly 46, 47, 48, 49, 50, 9, 51, 52. These deep networks are trained with the loss functions and deformation representations proposed in optimization based methods, but instead of iterating to find the optimal deformation, it is predicted directly. Such methods can be thought of as converting the homogenous and noisy intensity images into a feature image that is used to predict the deformation in a single step. Optimization methods study the family of deformations, their representations (elastostatics, viscous fluid, underlying Lie algebra, etc.) and how to optimize them, and learning focuses on automatic featurization of the intensity image that are conducive to registration.
Despite the extensive literature, Diffeomorphic Image Registration remains an active research area due to its high-dimensional solution space, ill-conditioned optimization, and non-Euclidean manifold of the transformation space. The significance of our work stems from the observation that these problems remain unaddressed by state-of-the-art optimization based registration methods, which typically use Gradient Descent 40, 53, 45 to optimize diffeomorphisms. In particular, first-order adaptive optimization methods are shown to speed up convergence in ill-conditioned optimization problems 54, 55, 56 without computing expensive second-order terms. Although first-order adaptive optimization methods have shown faster convergence to better local minima in fixed-dimensional Euclidean parameter spaces 55, 54, 56 (i.e. deep learning) and fixed low-dimensional non-Euclidean manifolds 57, 58, 59, these optimizers do not exist for diffeomorphic registration. This is because the size of the transform depends on the size of the image and changes over multi-scale optimization. We implement a novel multi-scale Adaptive Riemannanian optimization for Diffeomorphic Registration to mitigate the high-dimensional, ill-conditioned, non-Euclidean optimization problem. We introduce key technical contributions (Section 4) to avoid computing terms like the Riemannian Metric Tensor, and Parallel Transport of the optimization state, which are computationally expensive and not feasible for high-dimensional diffeomorphisms. To our knowledge, we are the first to implement a multi-scale Riemannian Adaptive Optimization algorithm for diffeomorphic registration. This leads to a state-of-the-art adaptive optimization algorithm for diffeomorphic registration(Figs. 2 and 3).
Our work also addresses the lack of scalability of existing registration algorithms. Existing optimization toolkits 60, 53, 61, 62 have prohibitively slow runtimes, which limits their applicability to hyperparameter studies for novel modalities or high-resolution images. Deep learning methods provide very fast runtimes but have steep compute and memory requirements, making them infeasible for high-resolution registration. Most deep learning methods perform registration on low-resolution image volumes 63 of size 160$\times$192$\times$224 voxels, which is much smaller than CT scans in EMPIRE10 10 (up to 420$\times$ 312 $\times$ 537 voxels) and RnR-ExM 64 (2048$\times$2048$\times$81 voxels) challenges. Most notably, for modalities like microscopy, existing methods 52, 62 either downsample the image volume by up to 64×64$\times$64 × or register image chunks independently. This aggressive downsampling or chunking leads to substantial loss of rich image features necessary for accurate registration. Our method can register these volumes at native resolution (Fig. 4), introducing a new benchmark for accurate and scalable image registration algorithms. This scalability also makes large-scale hyperparameter studies more computationally feasible (Figs. 6 and 5).
Our key contributions are as follows: first, we propose a novel multi-scale Adaptive Riemannan Optimization framework for diffeomorphisms. Our framework leverages mathematical correspondences to avoid expensive operations like the Riemannanian Metric Tensor and Parallel Transport which are needed for implementing first-order adaptive algorithms. This leads to a state-of-the art optimization algorithm that is accurate, fast and robust across various registration settings. Second, we accompany the method with a Python library that is easy to use and extend, and is packaged with optimizers for other transforms like rigid and affine transforms. Similar to existing toolkits 60, our method can compose transformations, avoiding resampling artifacts across transformations. This is designed to push the frontier of scalability in image registration algorithms. Our method scales in time, leading to up to a 3200 $\times$ speedup over existing state-of-the-art toolkits (Fig. 5) and scales in resolution, performing diffeomorphic registration on microscopy images at native resolution (Fig. 4). Our implementation is agnostic to modality, resolutions, and is not sensitive to hyperparameters, making it a versatile benchmark for diverse applications.
Figure 1: Overview of FireANTs and its features: (a) shows the modalities our method is tested on. We demonstrate results on in-vivo T1-weighted brain MRI, lung CT, and expansion microscopy volumes. FireANTs can optimize intensity images as well as binary masks (lung masks in CT) or entire anatomical label maps (brain MRI). (b) shows the technical contributions of FireANTs. We extend Adaptive Optimization to multi-scale Diffeomorphisms by first writing the Riemannian gradient update, and then avoiding parallel transport of the optimization state by leverging the interchangability of the Riemannian gradient at arbitrary transform $\varphi_{t}$ with the Riemannian gradient at $\varphi=\textbf{Id}$. For the Lie-algebra representation, the Gateaux derivative $\frac{\partial L}{\partial\varphi}$ is projected to $\frac{\partial L}{\partial v}$ using analytical backprop. Since the Lie algebra is a vector space, we use standard adaptive optimizers (see Section 4 for more details). (c) takes a closer look at multi-scale interpolation for diffeomorphisms represented as a warp field. Bicubic interpolation can introduce folding of the warp field at a finer resolution due to overshooting, but bilinear interpolation does not. Therefore, we use this for interpolating the warp field and the optimizer state. (d) shows the extensive experimental setup. Our method is orders of magnitude faster, has state-of-the-art performance on 3 challenges, is robust across hyperparameters and datasets, and is modular and easy to extend.
@@ 2 Results
Figure 2: FireANTs demonstrates state-of-the-art performance on Klein et al.challenge 5 : Registration quality is validated by measuring the average volume overlap measurements of all anatomical label maps between the fixed and warped label maps (see Section 4.1 for description of all metrics used). Our method outperforms state-of-the-art registration algorithms, including ANTs which was the top performing method in the challenge, and deep learning algorithms like VoxelMorph and SynthMorph. For deep learning baselines, appropriate preprocessing (intensity normalization, alignment, and resampling to 1mm isotropic) is performed to ensure a fair comparison, whereas no such preprocessing is required for optimization methods, including FireANTs. Boxplots show that the gains in performance are consistent across all four datasets, with the median overlap scores outperforming the third quartile of all other methods for IBSR and CUMC12 datasets. Results of per region overlap metrics are in Fig. S.1. For the overlap aggregation mentioned in 5, results are shown in Fig. S.2.
We validate the proposed features of our method using a comprehensive evaluation setup. First, we show that our proposed Riemannanian Adaptive Optimization leads to consistently better registration performance compared to state-of-the-art optimization algorithms that utilize Gradient Descent. This is shown on two challenges 5, 10 which are established community standards for evaluating registration algorithms, and present various challenges such as different modalities, anatomical regions and variations, voxel resolutions, anisotropy, and acquisition settings. Next, we demonstrate FireANTs’ scalability in resolution by performing deformation registration on high-resolution microscopy images 64 at native resolution, which are previously registered either at a significantly lower resolution or in chunks. We also show scalability in runtime by showing speedups of upto 3 orders of magnitude compared to existing SOTA algorithms. Finally, we show that FireANTs is robust to choice of registation hyperparameters, and its signifcant speedup allows for fast hyperparameter tuning, which is otherwise infeasible with existing algorithms.
2.1 Comparison on human brain MRI registration
@@
Analysis of functional and physiological data in neuroscience require different brain images to lie in the same coordinate space to establish correspondence across different brain regions. As such, registration algorithms are at the forefront of such analysis. Klein et al. 5 present a comprehensive, unbiased, and thorough evaluation of different registration algorithms on four MRI brain datasets, with Advanced Normalization Tools (ANTs) 60 being the top performing method overall. Four datasets are used in the challenge, with a total of 80 brains. The datasets have different voxel resolutions, scanners, preprocessing pipelines, and labeled anatomical regions. More details about the dataset and evaluation protocol are discussed in  Section 4.1. The challenge therefore evaluates robustness of registration algorithms across a wide variety of dataset attributes and anatomical alignment. We compare our method with two state-of-the-art optimization algorithms: ANTs and Symmetric Log Demons 61, and two deep learning algorithms: VoxelMorph 46 and SynthMorph 51 using their provided pretrained models. In addition to the proposed metrics in 5, we also propose alternate versions of the same metrics, but averaged over all the brain regions (see Section 4.1). For all the four datasets, we first fit an affine transformation from the moving image to the fixed image, followed by a diffeomorphic transform. Results for the brain datasets are shown in  Fig. 2 and  Fig. S.1.
@@
Our algorithm outperforms all baselines on all four datasets, with a monotonic improvement in all metrics evaluating the volume overlap of the fixed and warped label maps. The improvements are consistent in all datasets, with varying number and sizes of anatomical label maps. In the IBSR and CUMC12 datasets, the median target overlap of our method is better than the third-quantile of ANTs.  Fig. S.1 also highlights the improvement in label overlap per labeled brain region across all datasets. A small caveat with deep learning methods is that their performance is highly dependent on their domain gap between the training and test datasets. VoxelMorph is trained on the OASIS dataset, which has different image statistics compared to the four datasets, and consequently we see a performance drop. Moreover, VoxelMorph is sensitive to the anisotropy of the volumes, and all volumes have to be resampled to 1mm isotropic, and renormalized. A noticable performance drop is observed when the anisotropic volumes are fed into the network, which is undesirable as the trained model is essentially ‘locked’ to a single physical resolution - which limits the generalizability of the model to various modalities with different physical resolutions. For Demons, ANTs, and FireANTs (Ours), we do not perform any additional normalization or resampling. SynthMorph is more robust to the domain gap than VoxelMorph due to its training strategy with synthetic images, but still underperforms optimization baselines when their recommended hyperparameters are chosen.
2.2 Results on the EMPIRE10 lung CT challenge
(a) Comparison of fissure alignment error
@@ 2.2 Results on the EMPIRE10 lung CT challenge
(b) Singularity Errors in %
(c) Landmark distance across lung subregions
@@
(d) Fissure alignment error on top 10 algorithms in the challenge, averaged on all scan pairs
@@
Figure 3: FireANTs demonstrates state-of-the-art performance on EMPIRE10 lung registration challenge: (a): Lung fissure plates are an important anatomical landmark demarcating lobes within the lung. Fissure alignment errors (in %) denote the percentage of locations near the lung fissure plates that are on the wrong side of the fissure post-registration. Over all 30 scan pairs, our method performs 5$\times$ better than ANTs. (b): Singularity errors defined as percentage of voxels that have a non-diffeomorphic deformation. In the DARTEL baseline, singularities can be introduced for larger deformations due to numerical approximations of the integration. Even for the ANTs baseline, the solutions (deformations) returned are not entirely diffeomorphic. Our method shows much better fissure and landmark alignment ( Fig. 3(a,c),  Fig. S.4Fig. S.5) with guaranteed diffeomorphic transforms. (c): Landmark distance in mm for selected landmarks. Across different lung subregions, our method shows results at least at par with ANTs, with slightly better average and median results across all regions. (d): Shows the top 10 algorithms for average fissure alignment error in % in the EMPIRE10 challenge. Error metrics are taken from the evaluation server. Other methods perform well on one lung (MRF for right, ANTs for left) but comparatively poorly on the other lung, compared to our method showing both accurate and robustness to both the left and right lung.    X = First,    X = Second,    X = Third best result.
@@
Registration of thoracic CT data is one of the most common areas of research in the medical image registration community. The EMPIRE10 challenge  10 is an established benchmark challenge and it provides a platform for in-depth evaluation and fair comparison of available registration algorithms for this application. We discuss more details about the challenge in  Section 4.1. ANTs is, again one of the top performing methods in this challenge. Unlike the brain datasets, ground truth labels for fissure and landmarks are not provided for validation. Therefore, we rely on the evaluation metrics computed in the evaluation server. We compare our method with two powerful baselines (i) ANTs, which optimizes the diffeomorphism directly, and (ii) the DARTEL 53 formulation optimizing a stationary velocity field (SVF), where the diffeomorphism is obtained using an exponential map of the SVF. We first affinely align the binary lung masks of the moving and fixed images using Dice loss 65. This is followed by a diffeomorphic registration using the intensity images.

### v2

# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Matching

@@ Abstract
The paper proposes FireANTs, the first multi-scale Adaptive Riemannian Optimization algorithm for dense diffeomorphic image matching. One of the most critical and understudied aspects of diffeomorphic image matching algorithms are its highly ill-conditioned nature. We quantitatively capture the extent of ill-conditioning in a typical MRI matching task, motivating the need for an adaptive optimization algorithm for diffeomorphic matching. To this end, FireANTs generalizes the concept of momentum and adaptive estimates of the Hessian to mitigate this ill-conditioning in the non-Euclidean space of diffeomorphisms. Unlike common non-Euclidean manifolds, we also formalize considerations for multi-scale optimization of diffeomorphisms. Our rigorous mathematical results and operational contributions lead to a state-of-the-art dense matching algorithm that can be applied to generic image data with remarkable accuracy and robustness. We demonstrate consistent improvements in image matching performance across a spectrum of community-standard medical and biological correspondence matching challenges spanning a wide variety of image modalities, anatomies, resolutions, acquisition protocols, and preprocessing pipelines. This improvement is supplemented by from 300$\times$ up to 3200$\times$ speedup over existing state-of-the-art algorithms. For the first time, we perform diffeomorphic matching of sub-micron mouse cortex volumes at native resolution. Our fast implementation also enables hyperparameter studies that were intractable with existing correspondence matching algorithms.
Keywords: correspondence matching, deformable image matching, diffeomorphisms, optimization, non-Euclidean manifold, microscopy, neuroimaging
1 Introduction
The ability to identify and map corresponding elements across diverse datasets or perceptual inputs – known as correspondence matching – is fundamental to interpreting and interacting with the world. Correspondence matching between images is one of the longstanding fundamental problem in computer vision. Influential computer vision researcher Takeo Kanade famously once said that the three fundamental problems of computer vision are: “Correspondence, correspondence, correspondence” 1. Indeed, correspondence matching is fundamental and ubiquitous across various disciplines, manifesting in many forms including but not limited to stereo matching  2, structure from motion  3, 4, template matching  5, motion tracking  6, 7, shape correspondence  8, semantic correspondence  9, point cloud matching  10, optical flow  11, and deformable image matching  12. Solving these problems addresses the desiderata for a wide range of applications in computer vision, robotics, medical imaging, remote sensing, photogrammetry, geological and ecological sciences, cognitive sciences, human-computer interaction, self-driving among many other fields.
Correspondence matching is broadly divided into two categories: sparse and dense matching. Most sparse matching problems like stereo matching, structure from motion, and template matching involve finding a sparse set of salient features across images followed by matching them. In such cases, the transformation between images, surfaces, or point clouds is typically also parameterized with a small number of parameters, e.g., an affine transform, homography or a fundamental matrix. These methods are often robust to noise, occlusions, and salient features can be detected and matched efficiently via analytical closed forms. In contrast, dense matching is much harder because the entire image is considered for matching and cannot be reduced to a sparse set of salient features, and the transformation between images is typically parameterized with a large number of parameters, e.g., a dense deformation field. Moreover, dense matching is sensitive to local noise, and cannot be solved efficiently via analytical closed forms – necessitating iterative optimization methods 13, 14, 15, 16, 17, 18. Due to the dense and high-dimensional nature, these methods are often plagued with ill-posedness, curse of dimensionality, difficulty in optimization and scalability to high-resolution data.
In this work, we focus on dense deformable correspondence matching, which is the non-linear and local (hence deformable) alignment of two or more images into a common coordinate system. Dense deformable correspondence matching is a fundamental problem in computer vision 19, medical imaging 20, 21, 22, microscopy 23, 24, and remote sensing. Here, we focus on applications in biomedical and biological imaging. In the biomedical and biological sciences, deformable correspondence matching is also referred to as deformable registration. Within dense deformations, diffeomorphisms are of special interest as a family of deformations that are invertible transformations such that both the transform and its inverse are differentiable. This allows us to accurately model the deformation between images and also ensures that the deformation is physically plausible.
We address and tackle two fundamental problems in dense correspondence matching: ill-conditioning and scalability. The ill-conditioning arises due to the high-dimensional and heterogeneous nature of the dense matching optimization objective, that can be mitigated by adaptive optimization methods. Although standard adaptive optimization methods 25, 26 are shown to work in fixed Euclidean spaces, it is not obvious how to extend this formulation to the non-Euclidean space of diffeomorphisms. Fortunately, diffeomorphisms admit many interesting mathematical properties like being embedded in a Riemannanian manifold, having a Lie Group structure, and local geodesic formulations that can be exploited for adaptive optimization. We present a novel and mathematically rigorous framework for adaptive optimization of diffeomorphic matching Section 4.4. This is done by exploiting the group structure of diffeomorphisms to define a custom gradient descent algorithm, followed by adaptive optimization on this space. Second, we observe that most existing state-of-the-art methods are prohibitively slow for high-resolution data, which limits their applicability to rigorous hyperparameter studies, large-scale data, or high-resolution alignment at mesoscopic or microscopic resolutions. Our novel operational contributions lead to an algorithm that is 3200$\times$ faster than existing state-of-the-art toolkits (Fig. 4) and can perform diffeomorphic matching on modalities like microscopy images at native resolution (Fig. 3). This unprecedented speedup and scalability makes large-scale hyperparameter studies computationally feasible (Figs. 5 and 4). We package our contributions into a software toolkit called FireANTs, which is an open-source state-of-the-art toolkit for dense deformable correspondence matching. Our framework can allow a practitioner to perform interactive dense matching that is useful for aligning complex, multimodal, multi-channel or high-resolution images 27, 28 or to provide guidance to compensate for missing data (e.g., in microscopy imaging), wherein a typical dense image matching algorithm would fail to run interactively.
@@ 2 Results
FireANTs represents the next generation of algorithms superseding the widely established and successful adoption of the ANTs ecosystem spanning the gamut of biomedical and life sciences research. Accordingly, we rigorously evaluate FireANTs against established benchmarks to substantiate its efficacy and its viability as a robust registration algorithm across anatomies and modalities. Furthermore, we extend our analysis to encompass datasets with considerably higher image heterogeneity, namely deformable registration of a high-resolution expansion microscopy benchmark at native resolution. Finally, we demonstrate its unprecedented speedup and usability by showing speedups of up to three orders of magnitude and robustness to hyperparameters, subsequently opening avenues in the ability to perform large-scale hyperparameter studies and registration of exceedingly complex and larger datasets.
2.1 Experiment Setup
We briefly describe the significance, existing state-of-the-art and challenges associated with the chosen benchmarks to demonstrate the efficacy of FireANTs.
Klein et al. brain mapping challenge  29
Within the biomedical sciences, registration algorithms like ANTs have been predominantly used by the neuroimaging community by far. Klein et al. 29 in their landmark paper reported an extensive evaluation of fourteen state-of-the-art registration algorithms on four neuroimaging datasets. The four neuroimaging datasets (IBSR18, CUMC12, MGH10, LPBA40) comprise different whole-brain labelling protocols, eight different evaluation measures and three independent analysis methods. More details about the datasets are outlined in  Section A.1. ANTs was one of the top performing methods for this challenge, performing well robustly across all four datasets. Evaluating on this challenge is therefore imperative to establish FireANTs as an effective, versatile and robust algorithm for neuroimaging applications.
A natural way to evaluate whether two images are in a common coordinate frame is to evaluate the accuracy of overlap of gross morphological anatomical structures. The method considers measures of volume and surface overlap, volume similarity, and distance measures to evaluate the alignment of anatomical regions. Given a source label map ${S}_{r}$ and target label map ${T}_{r}$ and a cardinality operator |.||.|| . |, we consider the following overlap measures. The first measure ‘target overlap’, defined as the overlap between the source and target divided by the target. Target overlap is a measure of sensitivity, and the original evaluation 29 considers the aggregate total overlap as follows
$\small TO_{r}=\frac{|{S}_{r}\cap{T}_{r}|}{|{T}_{r}|},TO_{Klein}=\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}|{T}_{r}|}$ (1)
$\small TO=\frac{1}{N_{r}}\sum_{r}TO_{r}$ (2)
$MO_{r}=2\frac{|{S}_{r}\cap{T}_{r}|}{|{S}_{r}|+|{T}_{r}|},MO_{Klein}=2\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}(|{S}_{r}|+|{T}_{r}|)},MO=\frac{1}{N_{r}}\sum_{r}MO_{r}$ (3)
Klein et al.29 also propose a ‘Union Overlap’ metric which is a monotonic function of the Dice score. Therefore, we do not use this in our evaluation. To complement the above agreement measures, we also compute false negatives (FN), false positives (FP), and volume similarity (VS) coefficient for anatomical region $r$:
Similar to the overlap metrics, we compute the aggregates as in the original evaluation denoted by $FN_{Klein},FP_{Klein},VS_{Klein}$ and average over regions denoted simply by FN, FP, VS. This leads to a total of 10 aggregate metrics that we use to compare our method with 4 baselines - ANTs, Demons 18, VoxelMorph 30 and SynthMorph 31, representing established classical and deep learning registration algorithms.
EMPIRE10 challenge  22
Alignment of thoracic CT images, especially the lung and its internal structures is a challenging task, owing to the highly deformable nature of the lungs. Pulmonary registration has extensive clinical utility, such as aligning temporally distinct breath-hold scans to facilitate visual comparison and reduce errors. It also aids in modeling lung expansion by registering inspiration and expiration scans, and tracking disease progression by aligning breath-hold scans over time to enhance monitoring of airflow and pulmonary function. Murphy et al. propose the Evaluation of Methods for Pulmonary Image REgistration 2010 (EMPIRE10) challenge to provide a platform for a comprehensive evaluation and fair comparison of registration algorithms for the task of CT lung registration. The dataset consists of 30 scan pairs including inspiration-expiration scans, breathhold scans over time, scans from 4D data, ovine data, contrast-noncontrast, and artificially warped scan pairs. The challenge provides a broad range of data complexity, voxel sizes and image acquisition differences, and the challenge still continues as of today. The challenge only provides scan pairs and binary lung masks. All the other data (fissures and landmarks) are withheld for evaluation. All the scan pairs have varying spatial and physical resolutions, are acquired over a varying set of imaging configurations. This calls for a registration algorithm that is agnostic to any assumptions about anisotropy of image resolution, both physical and voxel. We use the evaluation provided by the challenge, and compare the fissure alignment, landmark alignment, and singularity of registration. More details about the evaluation can be found in  22. We compare our method with ANTs which performs direct gradient descent updates and DARTEL which optimizes a stationary velocity field using the metrics reported in the evaluation server.
RnR ExM mouse dataset  32
[…]
@@ 2.1 Experiment Setup
Figure 1: FireANTs demonstrates state-of-the-art performance on Klein et al.challenge 29: Following the evaluation setup of Klein et al. paper, we validate registration performance using the average volume overlap of all anatomical label maps between the fixed and warped label maps. We consider ANTs (the winner of the challenge), and Diffeomorphic Demons as state-of-the-art optimization algorithms, and Voxelmorph and Synthmorph as state-of-the-art unsupervised deep learning baselines. Evaluation is shown for five metrics with $\uparrow$ denoting a higher score is better, and $\downarrow$ signifying a lower score is better. For deep learning baselines, appropriate preprocessing (intensity normalization, alignment, and resampling to 1mm isotropic) is performed to ensure a fair comparison, whereas no such preprocessing is required for optimization methods, including FireANTs. FireANTs shows significant gains in performance that are consistent across all four datasets, with the median overlap scores outperforming the third quartile of all other methods for IBSR18 and CUMC12 datasets. Comparison of overlap metrics by specific anatomical regions are in Fig. 10. For the overlap aggregation mentioned in 29, results are shown in Fig. 11.
2.2 Results on state-of-the-art biomedical benchmarks
Klein et al. human brain MRI registration
@@ 2.2 Results on state-of-the-art biomedical benchmarks
Klein et al. 29 present a landmark challenge with comprehensive, unbiased, and thorough evaluation of different registration algorithms on four MRI brain datasets, with Advanced Normalization Tools (ANTs) 13 being the top performing method overall. We compare our method with two state-of-the-art optimization algorithms: ANTs - which won the original challenge, and Symmetric Log Demons 18, and two widely used deep learning algorithms: VoxelMorph 30 and SynthMorph 31 using their provided pretrained models. For all the four datasets, we first fit an affine transformation from the moving image to the fixed image, followed by a diffeomorphic transform. Results for the brain datasets are shown in  Fig. 1 and  Fig. 10.
@@
Our algorithm outperforms all baselines on all four datasets, with an improvement in all metrics evaluating the volume overlap of the fixed and warped label maps. The improvements are consistent in all datasets, with varying number and sizes of anatomical label maps. In the IBSR18 and CUMC12 datasets, the median target overlap of our method is better than the third-quantile of ANTs.  Fig. 10 also highlights the improvement in label overlap per labeled brain region across all datasets. A caveat with deep learning methods is that their performance is highly dependent on the domain gap between the training and test datasets. VoxelMorph is trained on the OASIS dataset, which has different image statistics compared to the four datasets, and consequently we see a performance drop. Moreover, VoxelMorph is sensitive to the anisotropy of the volumes, consequently all volumes are resampled to 1mm isotropic, and renormalized for VoxelMorph. A noticable performance drop is observed when the anisotropic volumes are fed into the network, which is undesirable as the trained model is essentially ‘locked’ to a single physical resolution - which limits the generalizability of the model to various modalities with different physical resolutions. For Demons, ANTs, and FireANTs (Ours), we do not perform any additional normalization or resampling. SynthMorph is more robust to the domain gap than VoxelMorph due to its training strategy with synthetic images, but still underperforms optimization baselines when their recommended hyperparameters are chosen.
(a) FireANTs achieves substantially lower inter-quartile range of fissure errors, defined as the percentage of marked pixels that are registered to points on the opposite side of the fissure boundary.
@@
(b) Singularity errors are defined as fraction of voxels that define a non-invertible deformation. Singularity quantifies the percentage of implausible deformations. FireANTs achieves zero percent singularity errors.
(c) Landmark distance is the Euclidean distance between well-dispersed landmark points between the fixed and warped images. FireANTs has a lower median and narrower interquartile range than baselines on five out of six subregions.
@@
(d) Fissure alignment error on top 10 algorithms in the challenge sorted by fissure alignment error, averaged on all scan pairs. FireANTs outperforms a wide array of baselines, including direct optimization (ANTs, ANTs+BSpline), neural networks (VIRNet), and explicit correlation volumes (MRF, Disco).
@@
Figure 2: FireANTs demonstrates state-of-the-art performance on EMPIRE10 lung registration challenge: (a): Lung fissure plates are an important anatomical landmark demarcating lobes within the lung. Fissure alignment errors (in %) denote the percentage of locations near the lung fissure plates that are on the wrong side of the fissure post-registration. Over all 30 scan pairs, our method performs 5$\times$ better than ANTs. (b): Singularity errors defined as percentage of voxels that have a non-diffeomorphic deformation, a proxy for physically implausible deformations. In the DARTEL baseline, singularities can be introduced for larger deformations due to numerical approximations of the integration. Even for ANTs, the solutions (deformations) returned are not entirely diffeomorphic. Our method shows much better fissure and landmark alignment (Fig. 2(a,c),  Fig. 13Fig. 14) with guaranteed diffeomorphic transforms. (c): Landmark distance in mm for selected landmarks. Across different lung subregions, our method shows results at least at par with ANTs, with slightly better average and median results across all regions. (d): Shows the top 10 algorithms for average fissure alignment error in % in the EMPIRE10 challenge. Error metrics are taken from the evaluation server. Other methods perform well on one lung (MRF for right, ANTs for left) but comparatively poorly on the other lung, compared to our method showing both accurate and robustness to both the left and right lung.    X = First,    X = Second,    X = Third best result.
EMPIRE10 lung CT challenge
@@
The EMPIRE10 challenge  22 is an established benchmark challenge and it provides a platform for in-depth evaluation and fair comparison of available registration algorithms for this application. We discuss more details about the challenge in  Section 2.1. The lung images represent volumes that are about 10$\times$ larger than the brain dataset, thereby presenting a significant challenge for deformable registration algorithms. ANTs is, again one of the top performing methods in this challenge. Unlike the brain datasets, ground truth labels for fissure and landmarks are not provided for validation. Therefore, we rely on the evaluation metrics computed in the evaluation server. We compare our method with two powerful baselines (i) ANTs, which optimizes the diffeomorphism directly, and (ii) the DARTEL 17 formulation optimizing a stationary velocity field (SVF), where the diffeomorphism is obtained using an exponential map of the SVF. We first affinely align the binary lung masks of the moving and fixed images using Dice loss 33. This is followed by a diffeomorphic registration using the intensity images.